# Spectra and pseudospectra in the evaluation of material stability

### v2

@@ 6.3 Structured pseudospectra
$A=\begin{bmatrix}\mathbf{0}&&&I&&&\mathbf{0}\\ A_{1}&&&\mathbf{0}&&&A_{2}\\ A_{3}&&&\mathbf{0}&&&A_{4}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;.$ (37)
In view of this structure, it may be appropriate to perturb only the blocks that are different from the identity and the null matrix. The notion of structured pseudospectrum enters thus into play [29], [35], [36].
@@
$\sigma_{\varepsilon}^{Str}(A)=\bigcup_{\|E\|<\varepsilon}\sigma(A+E)\;,$ (38)
@@
$E=\begin{bmatrix}\mathbf{0}&&&\mathbf{0}&&&\mathbf{0}\\ E_{1}&&&\mathbf{0}&&&E_{2}\\ E_{3}&&&\mathbf{0}&&&E_{4}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;.$ (39)
The complex pseudospectrum is computed adopting the close relation between the eigenvalues of the perturbed matrices and the resolvent norm. In the case of real perturbations, such a relation does not exist. Thus we proceed as follows: For any $\zeta\not\in\sigma(A)$, denote the resolvent by $\mathbf{R}(\zeta)=(\zeta I-A)^{-1}$. The 2-norm of the smallest real matrix $E$ that makes $\zeta$ an eigenvalue of $A+E$ is given by
$p(A,\zeta)\equiv\left(\inf_{\gamma\in(0,1]}s_{2}\left(\begin{bmatrix}\mathrm{Re}\,\mathbf{R}(\zeta)&&&-\gamma\mathrm{Im}\,\mathbf{R}(\zeta)\\ \gamma^{-1}\mathrm{Im}\,\mathbf{R}(\zeta)&&&\mathrm{Re}\,\mathbf{R}(\zeta)\end{bmatrix}\right)\right)^{-1}\;,$ (40)
where $s_{2}$ is the second-largest singular value. Thus, the 2-norm real structured $\varepsilon$-pseudospectrum of $A\in\mathbb{R}^{N\times N}$ is given by [36, Theorem 50.1]
$\sigma_{\varepsilon}^{\mathbb{R}}(A)=\{\zeta\in\mathbb{C}\;|\;p(A,\zeta)<\varepsilon\}\;.$ (41)
In the special case of quasicrystals treated here, however, the pseudospectrum analyses do not differ in essence from their structured counterparts. Figure 1 exemplifies such an aspect. Of course, it will not be always necessarily so.
@@ 7 Pseudospectrum in the quasicrystal dynamics: implications
We refer explicit computations to quasicrystals, as already anticipated. The software used is the one developed by Wright [40]. Data considered are as follows:
@@
Mass density is $\rho=5100\,\mathrm{kg\,m^{-3}}$ [39], [19]. Values of the coupling parameters $\chi$ and $\alpha$ as well as those of $k_{0}$ and $\varsigma$, which enter the self-action, are unknown. However, we know a theoretical range for the coupling parameter $\chi$, which is $[0,1.5]\,\mathrm{GPa}$, as established in [28]. We also take the friction coefficient $\varsigma$ of the form $\varsigma=e^{\mathfrak{N}}\,\mathrm{Pa\cdot s/m^{2}}$, with $\mathfrak{N}\in[15,19]$.
@@
In the following analysis of the pseudospectrum, perturbations $\varepsilon$ are of the order of Pa ($\varepsilon\approx 10^{-9}\chi$).
@@ 7.3 Microstructural self-action with purely dissipative character
In Figure 1 we compare complex and real $\varepsilon$-pseudospectra of $A$, the latter being a subset of the former, for different values of $\varepsilon$ in Pa, whose 10-based logarithm is displayed in the colorbar on the right of every snapshot. The material parameters are fixed to $\chi=0.5$ GPa and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$ ($\mathfrak{N}=19$). With these values, the departure from normality of $A$ given by the condition number $\kappa_{2}(V)$ is $\kappa(V)=6.49\cdot 10^{3}$.
(a) $\sigma_{\varepsilon}(A)$
(b) $\sigma_{\varepsilon}^{\mathbb{R}}(A)$
(c) $\sigma_{\varepsilon}^{Str}(A)$
Figure 1: Comparison between complex, real and structured $\varepsilon$-pseudospectrum of the matrix $A$ in (25) for $\chi=0.5$ GPa and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$. The figure on the right is made by the eigenvalues of 200 perturbation matrices of the form (39). In the figure (c) there is a zoom of the eigenvalue clouds: their values are, indeed, very close with each other.
Table 3 lists the eigenvalues of $A$ for $\alpha=0$ GPa, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$ as $\chi$ varies. When $\chi=1.7$ GPa the last two eigenvalues have positive real part: instability occurs. We find the same result by looking at the pseudospectrum in Figure 2 as $\chi$ varies, for $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, that is $\mathfrak{N}=19$, and $\alpha=0$ GPa. Once again instability is reached for $\chi>\sim 1.6$, so we recover numerically and from another path the theoretical result in [28] and we conclude that in this case the eigenvalues provide reliable information on the transition to instability.
@@
Figure 2: Dissipative self-action: real pseudospectrum in a neighborhood of the real eigenvalues of $A$ as $\chi$ varies.
Fixing $\chi=0.1$ GPa and $\alpha=0$ GPa, as $\mathfrak{N}$ varies we obtain results collected in Table 4 and Figure 3. Once again we show the pseudospectra next to the real eigenvalues of $A$, which are the ones that become unstable when $\chi>1.6$ GPa. As $\mathfrak{N}$ grows, the imaginary part of the complex eigenvalues of $A$ diminishes. These eigenvalues correspond to damped oscillations and describe asymptotically stable states: the origin is an attracting point in the phase space. As $\varsigma$ grows, the damping decreases; also the real part of the eigenvalues grows as $\mathfrak{N}$ increases. Correspondingly, we have non-oscillating solutions: the displacement tends exponentially to zero from the initial value, without oscillations. Finally, the presence of a dissipative phason self-action does not stabilize the matter in the (numerical) admissible range for the coupling parameter: for $\mathfrak{N}$ varying from $15$ to $20$, instability always occurs for $\chi\geq 1.69$ GPa. In particular, when $\chi<1.69$ GPa, the spectral radius of $A$ goes to zero as $\mathfrak{N}$ increases.
Table 4: Dissipative self-action: spectrum of $A$ for some values of $\mathfrak{N}$.
@@
Figure 3: Dissipative self-action: real pseudospectrum of $A$ as $\mathfrak{N}$ varies.
7.4 Simultaneous presence of conservative and dissipative components of z
When $\mathbf{z}=k_{0}\nu+\varsigma\dot{\nu}$, we consider the matrix in formula (28) with $\mathbb{K}_{i}$, $i=1,2,3$ given by formulas (29), (30), and (31).
Figure 4 and 5 show the complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa and $\chi=1.5$ GPa, respectively and for the same values of $\varepsilon$. Even if the matrix $A$ is genuinely non-normal, with $\kappa_{2}(V)=6.50\cdot 10^{3}$, its spectrum is stable under (generic) complex perturbations.
Figure 4: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa
Figure 5: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa
Table 5 collects eigenvalues of $A$ as $\chi$ varies, by fixing $\alpha=0$ GPa, $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, while Table 6 shows values of $\sigma(A)$ for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and $\alpha=0$ GPa, as $\chi$ varies. Increasing values of the conservative component of $\mathbf{z}$ have a stabilizing effect. If $\chi=3$ GPa, instability occurs when $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, while stability is recovered for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$.
@@ 7.4 Simultaneous presence of conservative and dissipative components of z
We fix $\chi=0.5$ GPa, $\alpha=0$ GPa, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and analyze the real psudospectrum of $A$ as $k_{0}$ varies. Results are in Figure 6: as $k_{0}$ grows, the three real eigenvalues of $A$ decrease; stability is assured.
Figure 6: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as a function of $k_{0}$.
If we allow $\varsigma$ to vary, by changing only $\mathfrak{N}$, and fixing $\chi=0.4$ GPa, $\alpha=0$ GPa, $k_{0}=0.09$ $\mathrm{GPa/m^{2}}$, as $\varsigma$ grows the real eigenvalues of $A$ tend toward 00 and, at the same time, the real part of the complex eigenvalues diminishes (see Figure 7). In this case, stability is assured for every value of $\mathfrak{N}$ considered.
Figure 7: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as $\mathfrak{N}$ varies.
Eventually, we allow $\chi$ to vary, keeping all the other parameters fixed. We set $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, $\alpha=0$ GPa, and $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$. As $\chi$ grows, the complex eigenvalues of $A$ move towards the half-plane with positive real part and for $\chi>\sim 1.8$ GPa there is a Hopf-type transition towards instability. By increasing $k_{0}$, instability occurs for higher values of $\chi$; for example, for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$ instability occurs when $\chi>3$ GPa, see table 6.
Figure 8: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as $\chi$ varies.
@@ 8 Concluding remarks
When we analyze the dynamics of complex bodies, especially when microstructural diffusion is present, the evolution of waves with time-dependent amplitude usually leads to (and recurrently does so) non-normal matrices. The parameterized spectral analysis, which appears natural when some constitutive coefficients are experimentally unknown or are affected by uncertainties, may not be sufficient to clearly reveal stability and possible transition to unstable states. Exploring pseudospectra gives more reliable estimates, as those obtained in the present paper in the case of quasicrystals, for which constitutive parameters pertaining to the phason self-action are still unknown; their possible range can be determined by an experimental campaign aimed at exploring the possible stability/instability behavior indicated theoretically by the above results. The same question emerges when we homogenize at continuum scale manufactured lattice-like structures at microscopic or mesoscopic scale (like in the case of printed metamaterials): a single lattice leads generically to a classical continuum scheme; multiple lattices imply often multi-field continuum schemes, those falling within the general model-building framework of the mechanics of complex materials (an offspring of this framework is what has been discussed here). In the last case, the occurrence of genuinely non-normal matrices in the material stability analysis seems to be more common. In all these case, recourse to pseudospectra is helpful if not necessary and requires the following steps:
@@
While in the special cases investigated in this paper we found that the pseudospectral analysis confirmed and validated the stability information obtained on the basis of the spectral analysis, in other cases this may not be necessarily so.
Acknowledgements. This research falls within the activities of the Theoretical Mechanics Research Group of the Mathematics Research Center E. De Giorgi of the Scuola Normale Superiore in Pisa. We acknowledge the support of GNFM-INDAM and GNCS-INDAM.

### v3

@@ 6.4 Structured pseudospectra
$A=\begin{bmatrix}\mathbf{0}&&&I&&&\mathbf{0}\\ A_{1}&&&\mathbf{0}&&&\frac{1}{\rho}A_{2}\\ \frac{1}{\varsigma}A_{2}&&&\mathbf{0}&&&A_{3}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;.$ (41)
where the $A_{i}$ blocks are symmetric $3\times 3$ matrices. In view of this structure, it may be appropriate to perturb only the blocks that are different from the identity and the null matrix. The notion of structured pseudospectrum enters thus into play [41], [47], [49].
@@
$\sigma_{\varepsilon}^{Str}(A)=\bigcup_{\|E\|_{2}<\varepsilon}\sigma(A+E)\;,$ (42)
@@
$E:=\begin{bmatrix}\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \frac{1}{\rho}E_{1}&\mathbf{0}&\frac{1}{\rho}E_{2}\\ \frac{1}{\varsigma}E_{2}&\mathbf{0}&\frac{1}{\varsigma}E_{3}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;,$ (43)
with $E_{i}\in Sym(\mathbb{R}^{3},\mathbb{R}^{3})$ and $\|E\|_{2}<\varepsilon$.
The structured $\varepsilon$-pseudospectrum is of course a subset of the complex $\varepsilon$-pseudospectrum, $\sigma_{\varepsilon}^{Str}(A)\subset\sigma_{\varepsilon}(A)$.
In our particular case the second and the third blocks in the matrices defined in equations (25) and (28) have different scales. To consider the circumstance that events described by $\nu$ occur at a scale different from the macroscopic one, we consider perturbations of the form (43); this way we are consistent with the physics discussed.
The complex pseudospectrum is computed considering the close relation between the eigenvalues of the perturbed matrices and the resolvent norm. In the case of real structured perturbations, such a relation does not exist. Thus, we proceed by computing the spectrum of $400$ perturbed matrices $A+E$, where $E$ has the structure given by (43).
6.5 On the values of the admissible perturbations
We look at perturbations of material parameters due to possible uncertainty in measurements of such parameters. Assume, for example, that the value of the coupling parameter $\chi$ is known up to an error of the order of $5\%$. Thus, we can write $\chi=\chi_{0}\pm\tilde{\chi}$, with $\tilde{\chi}=0.05\chi_{0}$ and $\chi_{0}$ the nominal value adopted. Since the matrix $A$ has linear dependence with respect to some of the material parameters (e.g. $\chi$ and the Lamè constants), the idea is to split $A$ into the sum of two matrices $A=A_{0}+\tilde{A}$, where $A_{0}$ is evaluated with the nominal values of the material parameters while $\tilde{A}$ is computed over the perturbations. Then we consider the 2-norm of the matrix $\tilde{A}$ as the maximum value of $\varepsilon$ physically meaningful for the computation of the structured $\varepsilon$-pseudospectrum. In the following, we say that a certain perturbation matrix $E$ with 2-norm equal to $\varepsilon$ corresponds to a perturbation of the order of $5\%$ on the values of the material parameters when the matrix $\tilde{A}$, evaluated with the perturbed values of the constitutive coefficients, has 2-norm equal to $\varepsilon$.
@@ 7 Pseudospectrum in the quasicrystal dynamics: implications
We refer explicit computations to quasicrystals, as already anticipated. The software used for the complex pseudospectrum is the one developed by Wright [54]. Data considered are as follows:
@@
Mass density is $\rho=5100\,\mathrm{kg\,m^{-3}}$ [53], [27]. Values of the coupling parameters $\chi$ and $\alpha$ as well as those of $k_{0}$ and $\varsigma$, which enter the self-action, are unknown. However, we know a theoretical range for the coupling parameter $\chi$, which is $[0,1.5]\,\mathrm{GPa}$, as established in [39]. We also take the friction coefficient $\varsigma$ of the form $\varsigma=e^{\phi}\,\mathrm{Pa\cdot s/m^{2}}$, with $\phi\in[15,19]$.
@@
The entries in $A$ are in practice always known only approximately. Among the sources of uncertainty in the matrix entries there are rounding errors due to finite precision computer arithmetic and those due to the experimental evaluation of physical parameters.
In the following analysis we consider both complex and structured pseudospectra with different purposes:
  • The complex pseudospectra $\sigma_{\varepsilon}(A)$ are a useful tool, e.g., for understanding if the eigenvalues are stable under perturbations due to operations in finite precision arithmetic. For this type of stability analysis, the values of $\varepsilon$ considered are thus of the order of machine precision.
  • The structured pseudospectra $\sigma^{Str}_{\varepsilon}(A)$ are used to understand if an uncertainty on the nominal values of the materials parameters may cause a transition towards instability before the one predicted by the eigenvalues. Thus, the order of the perturbations must have a physical meaning. In particular $\varepsilon$ must correspond to a certain variation (e.g., due to experimental uncertainty) of the values of the constitutive coefficients. Since, as shown in the results, the system becomes unstable for large values of the coupling parameter $\chi$, in the following we choose to adopt values of $\varepsilon$ corresponding to a variation of the order of $5\%$ of $\chi$. If any of the eigenvalues of perturbed matrices $A+E$, where $E$ is of the form (43), have positive real part, we conclude that the system must be regarded as unstable.
@@ 7.3 Microstructural self-action with purely dissipative character
For this matrix, the measures of departure from normality are $\kappa_{2}(V)=6.50\cdot 10^{3}$, $\text{dep}^{c}(A)=1$ (up to six digits) and $\text{dep}_{F}^{H}=4.58\cdot 10^{7}$. Thus, from the bound (37) we find
$1.52\cdot 10^{7}\leq{\text{dist}}(A,\mathfrak{K})\leq 4.48\cdot 10^{7}\;.$ (44)
These measures confirms that the matrix $A$ is genuinely non-normal. Also, they remain approximately constant for different values of $\chi$ and $\varsigma$ in the ranges considered in this work. Figure 1 shows the boundary of the 2-norm complex pseudospectra of $A$ for $\chi=0.1$ GPa and $\phi=19$. The ticks on the color-bar on the right correspond to the 10-based logarithm of $\varepsilon$, i.e. the contour correspond to $\varepsilon=10^{-0.75},10^{-1.5}$, etc. For $\varepsilon\leq 10^{-6}$, $\sigma_{\varepsilon}(A)$ is completely contained in the left half part of the complex plane. Even though $A$ is genuinely non-normal, its eigenvalues are stable under perturbations due to operations in finite precision. This is true for any value of the material parameters $\chi$ and $\varsigma$ considered.
Figure 1: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.1$ GPa and $\phi=19$. The dashed line is the imaginary axis.
Table 3 lists the eigenvalues of $A$ for $\alpha=0$ GPa, $\varsigma=0.178$ and $\mathrm{GPa\cdot s/m^{2}}$, as $\chi$ varies. When $\chi=1.7$ GPa the last two real eigenvalues are positive: instability occurs. We do not find the same result by looking at the structured pseudospectrum in Figure 2 as $\chi$ varies, for $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, that is, $\phi=19$ and $\alpha=0$ GPa. For $\chi\geq 0.4$ GPa, the structured $\varepsilon$-pseudospectra, for $\varepsilon$ corresponding to $5\%$ of $\chi$, crosses the imaginary axis: the real eigenvalues of some perturbed matrices $A+E$ are positive. The complex eigenvalues of $A$ are instead stable under such a type of perturbation, as shown in the clouds in the first plot of Figure 2. This behavior occurs for any value of the friction factor $\varsigma$ ($\phi$ varying from $15$ to $20$): when $\chi\geq 0.4$ GPa, the pseudospectra predict a transition towards instability of the system. We conclude that a dissipative component of the phason self-action does not change the (numerical) admissible range for the coupling parameter and, in this sense, it does not further stabilize the matter. Moreover, due to matrix non-normality, the analysis of the stability based on the spectrum $\sigma(A)$ does not correctly predict the transition towards instability.
@@
Figure 2: Dissipative self-action: structured pseudospectrum of $A$ as $\chi$ varies. Superposition of eigenvalues of $400$ matrices $A+E$, where $E$ is a matrix with random entries and the structure given in (43).
Fixing $\chi=0.1$ GPa and $\alpha=0$ GPa, as $\phi$ varies we obtain results collected in Table 4 and Figure 3. Once again in figure 3 we show the structured pseudospectra of $A$. Since $\chi<0.4$ GPa, instability is never reached. As $\phi$ grows, the imaginary part of the complex eigenvalues of $A$ diminishes. These eigenvalues correspond to damped oscillations and describe asymptotically stable states: the origin is an attracting point in the phase space. As $\varsigma$ grows, the damping decreases; also the real part of the eigenvalues grows as $\phi$ increases; in particular, when $\chi<0.4$ GPa, the spectral radius of $A$ goes to zero as $\phi$ increases. Correspondingly, we have non-oscillating solutions: the displacement tends exponentially to zero from the initial value, without oscillations.
Table 4: Dissipative self-action: spectrum of $A$ for some values of $\phi$.
@@
Figure 3: Dissipative self-action: structured pseudospectrum of $A$ as $\phi$ varies.
7.4 Simultaneous presence of conservative and dissipative components of $\mathbf{z}$
When $\mathbf{z}=k_{0}\nu+\varsigma\dot{\nu}$, we consider the matrix in formula (28) with $\mathbb{K}_{i}$, $i=1,2,3$ given by formulas (29), (30), and (31). It is genuinely non-normal, as confirmed by the measures of departure from normality: $\kappa_{2}(V)=6.50\cdot 10^{3}$, $\text{dep}^{c}(A)=1$ (up to six digits) and $\text{dep}_{F}^{H}=4.58\cdot 10^{7}$. We have
$1.52\cdot 10^{7}\leq{\text{dist}}(A,\mathfrak{K})\leq 4.48\cdot 10^{7}\;.$ (45)
Once again these measures remain almost constant for different values of $\chi$, $\varsigma$ and $k_{0}$. The departure from normality is the same as in the previous case. The spectrum of $A$ is stable under (generic) complex perturbations of the order of machine precision, as shown in figure 4. Again, each color corresponds to the boundary of $\sigma_{\varepsilon}(A)$ for different values of $\varepsilon$ whose 10-based logarithm are displayed in the color-bar on the right.
Figure 4: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=1.5$ GPa, $k_{0}=0.001$ GPa/m2, $\phi=17$. The dashed line is the imaginary axis.
Table 5 collects eigenvalues of $A$ as $\chi$ varies, by fixing $\alpha=0$ GPa, $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, while Table 6 shows values of $\sigma(A)$ for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and $\alpha=0$ GPa, as $\chi$ varies. Increasing the values of the conservative component of $\mathbf{z}$ has a stabilizing effect. If $\chi=3$ GPa, instability occurs when $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, while stability is recovered for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$.
@@ 7.4 Simultaneous presence of conservative and dissipative components of $\mathbf{z}$
We fix $\chi=0.5$ GPa, $\alpha=0$ GPa, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and analyze the structured pseudospectrum of $A$ as $k_{0}$ varies. Results are in Figure 5: as $k_{0}$ grows, the three real eigenvalues of $A$ decrease; stability is assured.
Figure 5: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: structured pseudospectrum of $A$ as a function of $k_{0}$.
If we allow $\varsigma$ to vary, by changing only $\phi$, and fixing $\chi=0.4$ GPa, $\alpha=0$ GPa, $k_{0}=0.09$ $\mathrm{GPa/m^{2}}$, as $\varsigma$ grows the real eigenvalues of $A$ tend toward 00 and, at the same time, the real part of the complex eigenvalues diminishes (see Figure 6). In this case, stability is assured for every value of $\phi$ considered.
Figure 6: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: structured pseudospectrum of $A$ as $\phi$ varies.
Eventually, we allow $\chi$ to vary, keeping all the other parameters fixed. We set $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, $\alpha=0$ GPa, and $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$. As $\chi$ grows, the three real eigenvalues of $A$ move towards the half-plane with positive real part and for $\chi>\sim 1.8$ GPa there is a Hopf-type transition towards instability. By increasing $k_{0}$, instability occurs for higher values of $\chi$; for example, for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$ instability occurs when $\chi>3$ GPa, see table 6.
However, the analysis of the structured pseudospectra shows that the real eigenvalues of $A$ are unstable under perturbations of the form (43) for $\varepsilon$ corresponding to $5\%$ of $\chi$. The structured $\varepsilon$-pseudospectrum of $A$ crosses the imaginary axis for values of $\chi$ smaller than the ones predicted by the eigenvalues, as shown in figures 7a and 8 for two different values of $k_{0}$. Instability is reached already for $\chi\geq 0.7$ GPa when $k_{0}=0.01$ GPa/m2 and for $\chi\geq 1.2$ GPa when $k_{0}=0.1$ GPa/m2. We emphasize that increasing the conservative component of the phason self-action expands the admissible range of $\chi$.
[…]
. Instability occurs for $\chi>0.6$ GPa.
[…]
@@ 8 Concluding remarks
When we analyze the dynamics of complex bodies, especially when microstructural diffusion is present, the evolution of waves with time-dependent amplitude may imply the emergence of non-normal matrices. The parameterized spectral analysis, which appears natural when some constitutive coefficients are experimentally unknown or are affected by uncertainties, may not be sufficient to clearly reveal stability and possible transition to unstable states. Exploring pseudospectra gives more reliable estimates, as those obtained in the present paper in the case of quasicrystals, for which constitutive parameters pertaining to the phason self-action are still unknown. The same question emerges when we homogenize at continuum scale manufactured lattice-like structures at microscopic or mesoscopic scale (like in the case of printed metamaterials): a single lattice leads generically to a classical continuum scheme; multiple lattices imply often multi-field continuum schemes, those falling within the general model-building framework of the mechanics of complex materials (an offspring of this framework is what has been discussed here), where the occurrence of genuinely non-normal matrices in the material stability analysis cannot be excluded a priori. In all these cases, recourse to pseudospectra is necessary for a reliable evalutation of the stability of the material. Indeed, as shown in the previous analyses, there are circumstances in which pseudospectra predict instability when, instead, the spectral analysis indicates stability.
The recourse to pseudospectra requires the following steps:
@@
The procedure is efficient and allows us to distinguish subtle behaviors that would otherwise be difficult to detect.
Acknowledgements. We thank the three reviewers for their kind evaluations and the suggestions. This research falls within the activities of the Theoretical Mechanics Research Group of the Mathematics Research Center E. De Giorgi of the Scuola Normale Superiore in Pisa. GNFM-INDAM and GNCS-INDAM are acknowledged.